# General effective reduction theory of integral polynomials of given non-zero discriminant and its applications

### v1

@@ 4.2. Explicit versions of theorems of Győry (1973) and Evertse and Győry (1991a)
    Denote by $\alpha_{1},\ldots,\alpha_{n}$ the zeros of $f$, and by $G$ the splitting field of $f$ over $\mathbb{Q}$. Then $[G:\mathbb{Q}]\leq n!$ and the absolute value $|D_{G}|$ of the discriminant of $G$ can be estimated from above by a constant $c_{9}(D)$. Here and below $c_{9},\ldots$ are effectively computable numbers depending only on $D$.
  2. (2)
    Putting $\Delta_{ij}:=\alpha_{i}-\alpha_{j}$ we have
@@
    which implies $|N_{G/\mathbb{Q}}\Delta_{ij}|\leq c_{10}(D)$. It follows that
    (4.3) $\Delta_{ij}=\delta_{ij}\varepsilon_{ij}\text{, where }H(\delta_{ij})\leq c_{11}(D)$
@@
  3. (3)
    The following identity plays a basic role in the proof:
@@
    Consider the graph, whose vertices are $\Delta_{ij}$ $(1\leq i\not=j\leq n)$ and whose edges are $[\Delta_{ij},\Delta_{ik}]$, $[\Delta_{ij},\Delta_{jk}]$ ($1\leq i\not=j\leq n$, $k\not=i,j$). This graph is obviously connected.
  4. (4)
    Equations (4.3) and (4.4) give rise to a ‘connected’ system of unit equations
@@
    where $\delta_{ijk}:=\delta_{ij}/\delta_{ik}$, $\tau_{ijk}:=\delta_{jk}/\delta_{ik}$ are non-zero elements of $G$ with heights effectively bounded above in terms of $|D|$ only, and $\varepsilon_{ijk}:=\varepsilon_{ij}/\varepsilon_{ik}$, $\nu_{ijk}:=\varepsilon_{jk}/\varepsilon_{ik}$ are unknown units in the ring of integers of $G$.
  5. (5)
    Represent $\varepsilon_{ijk}$ as
    $\varepsilon_{ijk}=\zeta_{ijk}\varrho_{1}^{a_{ijk,1}}\cdots\varrho_{r}^{a_{ijk},r},$
    and similarly $\nu_{ijk}$, where $\zeta_{ijk}$ is a root of unity, $\varrho_{1},\ldots,\varrho_{r}$ a fundamental system of units in $G$ with heights effectively bounded above in terms of $|D|$ and, by Dirichlet theorem $r\leq n!-1$.
  6. (6)
    Applying Baker’s method to (4.5), we get effective upper bounds for $|a_{ijk,1}|,\ldots,|a_{ijk},r|$ that depend only on $|D|$. In view of (4.3), this leads to upper bounds for the heights of the quotients $\Delta_{ij}/\Delta_{ik}=\delta_{ijk}\varepsilon_{ijk}$ for each triple $\{i,j,k\}\subset\{1,\ldots,n\}$, depending on $G,n$ and $|D|$, and so eventually only on $|D|$, and likewise for $\Delta_{jk}/\Delta_{ik}$.
    Remark: in Győry (1974), this was the first application of Baker’s method to general unit equations of the form (4.5) with explicit bounds.
  7. (7)
    Using the connectedness of the unit equations involved, this yields effective upper bounds for the height of $\Delta_{ij}$ for every $i,j$, depending only on $|D|$. Indeed, one first obtains an upper bound for the height of any quotient $\Delta_{ij}/\Delta_{kl}$ via
@@
  8. (8)
    Adding the differences $\Delta_{ij}=\alpha_{i}-\alpha_{j}$ for fixed $i$ and for $j=1,\ldots,n$, using the fact that $\alpha_{1}+\cdots+\alpha_{n}\in\mathbb{Z}$, putting $\alpha_{1}+\cdots+\alpha_{n}=na+a^{\prime}$ with $a,a^{\prime}\in\mathbb{Z}$, $0\leq a^{\prime}<n$, and writing
@@
    we have that $g(X)=f(X+a)\in\mathbb{Z}[X]$ and that the height of $g$ has an effective upper bound depending only on $|D|$. ∎
We note that for cubic and quartic monic polynomials $f\in\mathbb{Z}[X]$ of given non-zero discriminant Klaska (2021, 2022) devises another approach for proving Corollary 4.3 via the theory of integral points on elliptic curves.
@@
Take an integral polynomial $f\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D\neq 0$. In view of Theorems 2.1, 2.2 and 2.4 we may assume that $n\geq 4$. The discriminant of the splitting field of $f$ can be estimated from above in terms of $D$, and by the Hermite–Minkowski Theorem, this leaves only a finite, effectively determinable collection of possible splitting fields for $f$. So we may restrict ourselves to polynomials $f$ with given splitting field $G$ and ring of integers $\mathcal{O}_{G}$.
@@
(4.6) $f=\prod_{i=1}^{n}(\alpha_{i}X-\beta_{i})\,\,\text{over }\overline{\mathbb{Q}},$
such that the number of linear factors with real coefficients is maximal, and the factors with complex coefficients fall apart into complex conjugate pairs. After multiplying $f$ by a small positive rational integer, which can be effectively bounded in terms of $G$, hence in terms of $D$ and which is negligible compared with the other estimates arising from the application of Baker’s method, we may assume that $f$ has such a factorization with $\alpha_{i},\beta_{i}\in\mathcal{O}_{G}$ for $i=1,\ldots,n$. Put
@@
  1. (1)
    We start with a small variation on the reduction theory of Hermite (1848, 1851) and Julia (1917). Let ${\bf t}=(t_{1},\ldots,t_{n})$ be a tuple of positive reals such that $t_{i}=t_{j}$ for each pair $(i,j)$ such that $\alpha_{i},\beta_{i}$ are the complex conjugates of $\alpha_{j},\beta_{j}$. Consider the positive definite quadratic form
@@
  2. (2)
    For any quadruple $i,j,k,l$ of distinct indices we have the identity
@@
    Notice that all terms $\Delta_{ij}$ are in $\mathcal{O}_{G}$ and divide $D$. Hence $|N_{G/\mathbb{Q}}(\Delta_{ij})|\leq|D|^{[G:\mathbb{Q}]}$ for all $i,j$ where $[G:\mathbb{Q}]\leq n!$. As above in Section 4.3, we can express each term $\Delta_{ij}$ as a product of an element of height effectively bounded in terms of $n,D$ and a unit from $\mathcal{O}_{G}$. By substituting this into the identities (4.7) we obtain homogeneous unit equations in three terms. Dividing (4.7) by $\Delta_{ik}\Delta_{jl}$ we get unit equations like in (4.5) above, and using the theorem of Győry (1974) we obtain effective upper bounds for the heights of the quotients $\Delta_{ij}\Delta_{kl}/\Delta_{ik}\Delta_{jl}$.
  3. (3)
    To obtain an effective upper bound for the height of $g$ in terms of $D$, it suffices to effectively estimate the quantities $M$ and $R$ from Proposition 4.9 from above in terms of $D$, for a suitable choice of the $t_{i}$. For the $t_{i}$ we choose

### v2

@@ 4.4. Unit equations and $S$-unit equations
1. Denote by $\alpha_{1},\ldots,\alpha_{n}$ the zeros of $f$, and by $G$ the splitting field of $f$ over $\mathbb{Q}$. Then $G:\mathbb{Q}]\leq n!$ and the absolute value $|D_{G}|$ of the discriminant of $G$ can be estimated from above by a constant $c_{11}(n,|D|)$. Here and below $c_{11},\ldots$ are effectively computable numbers depending only on $n$ and $|D|$.
2. Putting $\Delta_{ij}:=\alpha_{i}-\alpha_{j}$ we have
@@
which implies $|N_{G/\mathbb{Q}}\Delta_{ij}|\leq c_{12}(n,|D|)$. It follows that
(4.8) $\Delta_{ij}=\delta_{ij}\varepsilon_{ij}\text{, where }H(\delta_{ij})\leq c_{13}(n,|D|)$
@@
3. The following identity plays a basic role in the proof:
@@
Consider the graph, whose vertices are $\Delta_{ij}$ $(1\leq i\not=j\leq n)$ and whose edges are $\Delta_{ij},\Delta_{ik}]$, $\Delta_{ij},\Delta_{jk}]$ ($1\leq i\not=j\leq n$, $k\not=i,j$). This graph is obviously connected.
4. Equations (4.8) and (4.9) give rise to a ‘connected’ system of unit equations
@@
where $\delta_{ijk}:=\delta_{ij}/\delta_{ik}$, $\tau_{ijk}:=\delta_{jk}/\delta_{ik}$ are non-zero elements of $G$ with heights effectively bounded above in terms of $n$ and $|D|$ only, and $\varepsilon_{ijk}:=\varepsilon_{ij}/\varepsilon_{ik}$, $\nu_{ijk}:=\varepsilon_{jk}/\varepsilon_{ik}$ are unknown units in the ring of integers of $G$.
5. Applying Theorem 4.9, together with the Remark following it, we get upper bounds for the heights of the quotients $\Delta_{ij}/\Delta_{ik}=\delta_{ijk}\varepsilon_{ijk}$ for each triple $\{i,j,k\}\subset\{1,\ldots,n\}$, depending on $G,n$ and $|D|$, and so eventually only on $n$ and $|D|$, and likewise for $\Delta_{jk}/\Delta_{ik}$.
6. Using the connectedness of the unit equations involved, this yields effective upper bounds for the height of $\Delta_{ij}$ for every $i,j$, depending only on $n$ and $|D|$. Indeed, one first obtains an upper bound for the height of any quotient $\Delta_{ij}/\Delta_{kl}$ via
@@
7. Adding the differences $\Delta_{ij}=\alpha_{i}-\alpha_{j}$ for fixed $i$ and for $j=1,\ldots,n$, using the fact that $\alpha_{1}+\cdots+\alpha_{n}\in\mathbb{Z}$, putting $\alpha_{1}+\cdots+\alpha_{n}=na+a^{\prime}$ with $a,a^{\prime}\in\mathbb{Z}$, $0\leq a^{\prime}<n$, and writing
@@
we have that $g(X)=f(X+a)\in\mathbb{Z}[X]$ and that the height of $g$ has an effective upper bound depending only on $n$ and $|D|$. ∎
Remark.
We note that for cubic and quartic monic polynomials $f\in\mathbb{Z}[X]$ of given non-zero discriminant, Klaska (2021, 2022) devised another approach for proving Corollary 4.3 via the theory of integral points on elliptic curves.
@@
Take an integral polynomial $f\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D\neq 0$. In view of Theorems 2.1, 2.2 and 2.4 we may assume that $n\geq 4$. The absolute value of the discriminant of the splitting field of $f$ can be estimated from above in terms of $|D|$, and by the Hermite-Minkowski Theorem, this leaves only a finite, effectively determinable collection of possible splitting fields for $f$. So we may restrict ourselves to polynomials $f$ with given splitting field $G$ and ring of integers $\mathcal{O}_{G}$.
@@
[…]
such that the number of linear factors with real coefficients is maximal, and the factors with complex coefficients fall apart into complex conjugate pairs. After multiplying $f$ by a small positive rational integer, which can be effectively bounded in terms of $G$, hence in terms of $n$ and $|D|$ and which is negligible compared with the other estimates arising from the application of Baker’s method, we may assume that $f$ has such a factorization with $\alpha_{i},\beta_{i}\in\mathcal{O}_{G}$ for $i=1,\ldots,n$. Put
@@
1. We start with a small variation on the reduction theory of Hermite (1848, 1851) and Julia (1917). Let ${\bf t}=(t_{1},\ldots,t_{n})$ be a tuple of positive reals such that $t_{i}=t_{j}$ for each pair $(i,j)$ such that $\alpha_{i},\beta_{i}$ are the complex conjugates of $\alpha_{j},\beta_{j}$. Consider the positive definite quadratic form
@@
2. For any quadruple $i,j,k,l$ of distinct indices we have the identity
@@
Notice that all terms $\Delta_{ij}$ are in $\mathcal{O}_{G}$ and divide $D$. Hence $|N_{G/\mathbb{Q}}(\Delta_{ij})|\leq|D|^{[G:\mathbb{Q}]}$ for all $i,j$ where $G:\mathbb{Q}]\leq n!$. As above in Section 4.4, we can express each term $\Delta_{ij}$ as a product of an element of height effectively bounded in terms of $n,D$ and a unit from $\mathcal{O}_{G}$. By substituting this into the identities (4.12) we obtain homogeneous unit equations in three terms. Dividing (4.12) by $\Delta_{ik}\Delta_{jl}$ we get unit equations like in (4.10) above, and using Theorem 4.9 we obtain effective upper bounds for the heights of the quotients $\Delta_{ij}\Delta_{kl}/\Delta_{ik}\Delta_{jl}$.
3. To obtain an effective upper bound for the height of $g$ in terms of $n$ and $|D|$, it suffices to effectively estimate the quantities $M$ and $R$ from Proposition 4.13 from above in terms of $n$ and $|D|$, for a suitable choice of the $t_{i}$. For the $t_{i}$ we choose